OpenStudy (anonymous):

What is an equation of the line in slope-intercept form? m=2/7 and y intercept is (0-12)

OpenStudy (anonymous):

what is y intercept??

OpenStudy (anonymous):

is this (0,12)

OpenStudy (anonymous):

the y intercept is (0,-12)

OpenStudy (anonymous):

we have equation \[m=y _{2}-y _{1}/x _{2}-x_{1}\]

OpenStudy (anonymous):

we will use the equation as 2/7=y-(-12)/x-0

OpenStudy (anonymous):

solving this equation we will get \[\frac{ 2 }{ 7 }=\frac{ y+12 }{ x }\]

OpenStudy (anonymous):

now by cross multiplication you can solve this equation

OpenStudy (anonymous):

7*(y+12)=2*x

OpenStudy (dobby1):

y=mx+b is slope intercept form

OpenStudy (dobby1):

\[y=\frac{ 2 }{ 7 }x-12\]

OpenStudy (dobby1):

@mrivera3113 can i have a medal?

OpenStudy (anonymous):

A metal?

OpenStudy (dobby1):

yeah click best response next to my name